OpenStudy (anonymous):

How do i solve this by completing the square? 2x^2 + 20x - 190 = 0

OpenStudy (anonymous):

2(x^2 + 10x) - 190 = 0 2(x^2 + 10x + 25) - 190 - 50 = 0 2(x + 5)^2 - 240 = 0 2(x + 5)^2 = 240 (x + 5)^2 = 120 Can you take it from here?

OpenStudy (anonymous):

im a little confused as to how you got from step one to get the (x+5)^2 part

OpenStudy (anonymous):

@tcarroll010